OpenStudy (anonymous):

can someone help explain how you find the limit of this? thanks lim as x approaches 1+ ((sqrt 2x)(x-1))/(abs x-1) then lim as x approaches 1-

OpenStudy (anonymous):

In case that ^^ wasn't clear enough, I've attached an image of the problem

OpenStudy (dumbcow):

for part A when approaching from the right \[|x-1| = x-1\] so that they cancel, leaving a limit of sqrt(2) for part B, x being less than 1 you have \[|x-1| = -(x-1)\] after cancelling you left with limit of -sqrt(2)

OpenStudy (anonymous):

Well, that was embarrassingly simple. Thank you!